OpenStudy (anonymous):

Solve using the square root method. Round the solution to the nearest tenth. x2 - 20 = 0 A. {-10.0, 10.0} B. {-5.5, 5.5} C. {-3.5, 3.5} D. {-4.5, 4.5}

OpenStudy (saifoo.khan):

D

OpenStudy (anonymous):

x = sqrt{20} x = 2sqrt{5} D ?

OpenStudy (anonymous):

D

OpenStudy (across):

All you have to do is find\[\sqrt{20}.\]Well,\[5^2=25\]is too large and\[4^2=16\]is too small. Do you think it could perhaps be in between 4 and 5?